A ball is thrown vertically upwards at 29.4 m/s. For its complete trip (up and back down to the starting position), its average speed is:

a. not given.

b. 9.80 m/s.

c. 14.7 m/s.

d. 4.90 m/s.
Answer #1

Given that

Initial velocity of ball,u = 29.4 m/s

We know that

acceleration due to gravity = 9.8

v at maximum height = 0

Then for maximum height , {v}^{2} = {u}^{2} + 2ah

0 = 29.42 + 2 (-9.8) h

0 = 864.36 - 19.6 h

h = 44.1 m

Now for time
v = u + at
=> 0 = 29.4 + (-9.8) t

=> t = 3 s

now as total distance = 2 h = 2 (44.1) = 88.2 m
and total time = 2 t = 2(3) = 6 sec

then average velocity = Total distance/total time = 88.2/6 = 14.7 m/s

Average velocity = 14.7 m/s